# Collision-Free Navigation of Wheeled Mobile Robots: An Integrated Path Planning and Tube-Following Control Approach

### v1

@@ III-C Adaptive tube-following control
where $\rho>0$ is a user-defined constant. At each time instance, the actual position $\mathbf{x}$ of the virtual control point $\bm{P}$ is allowed to stay within a 1-sphere of radius $\rho$ and center $\mathbf{x}_{d}(t)$. Unifying all such spheres along $t$ forms a tube around $\mathbf{x}_{d}(t)$, as shown in Fig. 5. Since a safety margin is considered in the planning module, if the tube radius is taken no larger than the size $\epsilon$ of the safety margin and (24) holds for all $t\geq 0$, then the actual trajectory $\mathbf{x}(t)$ will remain within the predefined tube without colliding with any actual obstacles, that is, $\mathbf{x}\in\mathcal{X}$.
Figure 5: Illustration of the predefined tube.
In the following, an adaptive tube-following controller is developed to achieve high-precision trajectory tracking, while ensuring that $\mathbf{x}(t)$ evolves within the tube defined by (24). To this end, we define a transformed error
@@
Inspecting (25) reveals that $10\leq\xi(t)<$ 1 is equivalent to (24), and $\xi(t)=0$ only when $\mathbf{x}_{e}(t)=\mathbf{0}$. Therefore, the tube-following control problem boils down to achieving $\lim_{t\to\infty}\xi(t)=0$, while ensuring $10\leq\xi(t)<$ 1, $\forall t\geq 0$, via a properly-designed controller. To achieve this, the following logarithmic barrier function is considered
$L=\dfrac{1}{2}\ln\dfrac{1}{1-\xi}.$ (27)
Let $\mathbf{z}:=\mathbf{x}_{e}/(\rho^{2}(1-\xi))$. Then, evaluating the time derivative of $L$ along (26) yields
$\dot{L}=$ $~{}\mathbf{z}^{\top}[\mathbf{R}(\theta)(\mathbf{u}+\mathbf{u}_{d})-\bm{\tau}_{d}]$
$\leq$ $~{}\mathbf{z}^{\top}(\mathbf{R}(\theta)\mathbf{u}-\bm{\tau}_{d})+d\|\mathbf{z}\|,$ (28)
wherein we have used the fact that $\|\mathbf{R}(\theta)\|=1$. The adaptive control law is designed as
$\mathbf{u}=\mathbf{R}^{-1}(\theta)(-k\mathbf{x}_{e}+\bm{\tau}_{d}-\bm{\varpi}),$ (29)
@@
$\bm{\varpi}:=\dfrac{\hat{d}^{2}\mathbf{z}}{\sqrt{\hat{d}^{2}\|\mathbf{z}\|^{2}+\varphi^{2}}},$ (30)
where $\hat{d}$ denotes the estimate of $d$ (see Assumption 1 for its definition) and is updated by a projection-based adaptive law. Suppose $d\in\Omega:=\{\nu\in\mathbb{R}\mid 0\leq\nu\leq d_{m}\}$ and let $\Omega_{\delta}:=\{\nu\in\mathbb{R}\mid 0\leq\nu\leq d_{m}+\delta\}$, where $d_{m}>0$ and $\delta>0$ are known constants. Inspired by [29], the adaptive law is derived as
$\dot{\hat{d}}=\text{Proj}(\hat{d},\Phi),~{}\Phi:=\|\mathbf{z}\|-\gamma\hat{d},$ (31)
with
$\text{Proj}(\hat{d},\Phi)=\left\{\begin{split}&\eta\Phi,&{\rm if}~{}&1)~{}0\leq\hat{d}<d_{m}~{}\text{or}\\ &&~{}&2)~{}\hat{d}\geq d_{m}~{}\text{and}~{}\Phi\leq 0,\\ &\eta\breve{\Phi},&{\rm if}~{}&3)~{}\hat{d}\geq d_{m}~{}\text{and}~{}\Phi>0,\end{split}\right.$ (32)
where $\eta,\,\gamma>0$ are design constants and
$\breve{\Phi}=\left(1-\dfrac{\hat{d}-d_{m}}{\delta}\right)\Phi.$ (33)
Note that the projection operator $\text{Proj}(\hat{d},\Phi)$ is locally Lipschitz in $(\hat{d},\Phi)$ [29]. Choosing the initial condition of (31) as $\hat{d}(0)\in\Omega_{\delta}$, we will now demonstrate that
$\hat{d}(0)\in\Omega_{\delta}\Rightarrow\hat{d}(t)\in\Omega_{\delta},~{}\forall t\geq 0.$ (34)
In Cases 1 and 2, the adaptive law (31) reduces to
$\dot{\hat{d}}=\eta(\|\mathbf{z}\|-\gamma\hat{d}).$ (35)
Solving (35) yields
[…]
where $t_{0}\geq 0$ marks the beginning of each time interval during which Cases 1 and 2 hold. Moreover, in these two cases, (34) holds as a result of the given conditions. In Case 3, however, (31) becomes
$\dot{\hat{d}}=\dfrac{\delta+d_{m}-\hat{d}}{\delta}\Phi,$ (37)
which implies that $\hat{d}$ decreases when $\hat{d}>\delta+d_{m}$ and increases if $\hat{d}<\delta+d_{m}$ until it reaches $\delta+d_{m}$ such that $\dot{\hat{d}}=0$. Therefore, the adaptive law (31) ensures that (34) holds. Let us define the parameter estimation error as $\tilde{d}:=\hat{d}-d$.
@@
Consider the WMR kinematics (3) under Assumption 1 and suppose that $d\in\Omega$. Given the initial position $\mathbf{x}(0)$ satisfying (25), then the adaptive controller (29), together with the projection-based update law (31), can steer $\mathbf{x}$ to track the reference trajectory $\mathbf{x}_{d}$ given by (23) along with the predefined tube (25), while ensuring uniform ultimate boundedness of the position error $\mathbf{x}_{e}$ and the estimation error $\tilde{d}$.
Proof. Consider the Lyapunov function candidate (LFC):
$V=L+\dfrac{1}{2\eta}\tilde{d}^{2}.$ (38)
Taking the derivative of $V$ and using (III-C) and (29) lead to
$\dot{V}\leq-k\mathbf{z}^{\top}\mathbf{x}_{e}-\mathbf{z}^{\top}\bm{\varpi}+\hat{d}\|\mathbf{z}\|+\dfrac{1}{\eta}\tilde{d}(\dot{\hat{d}}-\eta\|\mathbf{z}\|).$ (39)
Invoking the inequality $|\nu|\leq\nu^{2}/\sqrt{\nu^{2}+\varphi^{2}}+\varphi$ for any $\nu\in\mathbb{R}$ and $\varphi>0$, one gets $-\mathbf{z}^{\top}\bm{\varpi}+\hat{d}\|\mathbf{z}\|\leq\varphi$. In addition, noting (26), (III-C), and Lemma 1, it follows that $-k\mathbf{z}^{\top}\mathbf{x}_{e}=-k\xi/(1-\xi)\leq-2kL$ for all $10\leq\xi<$ 1. As a result, (39) reduces to
$\dot{V}\leq-2kL+\varphi+\dfrac{1}{\eta}\tilde{d}(\dot{\hat{d}}-\eta\|\mathbf{z}\|),$ (40)
on the set $10\leq\xi<$ 1. To proceed, we consider three cases in (32). For Cases 1 and 2, it follows from (35) that
$\tilde{d}(\dot{\hat{d}}-\eta\|\mathbf{z}\|)=-\eta\gamma\tilde{d}\hat{d}.$ (41)
While for Case 3, we have
$\tilde{d}(\dot{\hat{d}}-\eta\|\mathbf{z}\|)=-\eta\gamma\tilde{d}\hat{d}-\dfrac{\eta\tilde{d}(\hat{d}-d_{m})\Phi}{\delta}\leq-\eta\gamma\tilde{d}\hat{d},$ (42)
which is true because $\tilde{d}\geq 0$, $\hat{d}-d_{m}\geq 0$, and $\Phi>0$ according to the conditions specified in Case 3. In view of (41) and (42), one can easily get
$\dot{V}$ $\leq-2kL-\gamma\tilde{d}\hat{d}+\varphi$
$\leq-2kL-\dfrac{\gamma}{2}\tilde{d}^{2}+c$
$\leq-aV+c,$ (43)
where $a:=\min\{2k,\eta\gamma\}>0$ and $c:=\gamma d^{2}/2+\varphi>0$. Integrating both sides of (III-C) over $[0,\infty)$ yields
$V(t)\leq\left(V(0)-\dfrac{c}{a}\right)e^{-at}+\dfrac{c}{a},~{}\forall t\geq 0,$ (44)
showing that $V\in\mathcal{L}_{\infty}$ and hence $L\in\mathcal{L}_{\infty}$. The later implies that $10\leq\xi(t)<$ 1 for all $t\geq 0$. As a consequence, $\mathbf{x}$ keeps within the predefined tube (25). As $L(t)\leq V(t)$, it follows from (25) and (27) that $\|\mathbf{x}_{e}(t)\|\leq\rho\sqrt{1-e^{-2V(t)}}$. In addition, $\tilde{d}^{2}/(2\eta)\leq V(t)$ leads to $|\tilde{d}(t)|\leq\sqrt{2\eta V(t)}$. As shown in (44), $V(t)$ is eventually bounded by $c/a$, whereby one can get
$\mathop{\lim\sup}\limits_{t\to\infty}\|\mathbf{x}_{e}(t)\|\leq\rho\sqrt{1-e^{-2c/a}},~{}\mathop{\lim\sup}\limits_{t\to\infty}|\tilde{d}(t)|\leq\sqrt{2\eta c/a}.$
@@
Based on the facts that $\|\mathbf{R}^{-1}(\theta)\|=1/\ell$ ($\ell<1$), $\|\bm{\tau}_{d}\|\leq\|\bm{\kappa}_{0}(\mathbf{x}_{d})\|\leq\alpha$ (noting (13), (14), and $\|\bm{\Pi}(\mathbf{x}_{d})\|=1$), and $\|\bm{\varpi}\|\leq\hat{d}\leq d_{m}+\sigma$ (noting (34)), it can be easily verified that the adaptive controller (29) is bounded and its bound can be explicitly expressed as
$\|\mathbf{u}\|\leq\dfrac{1}{\ell}(k\rho+\alpha+d_{m}+\delta).$ (45)
This implies that the input constraint $\|\mathbf{u}\|\leq u_{m}$ can be met by judiciously choosing the design parameters such that $k\rho+\alpha+d_{m}+\delta\leq\ell u_{m}$. However, it is noted that achieving smaller convergence bounds for $\mathbf{x}_{e}$ and $\tilde{d}$ requires a larger value of $k$. Thus, a compromise should be reached according to the robot actuation capability and performance requirement.
Remark 5.
As defined in (1), the kinematics of $\theta$ is
$\dot{\theta}=\omega+u_{d2},$ (46)
[…]
@@ IV Simulation Results
In this section, we demonstrate the proposed control method in a $6.4\,\text{m}\times 3.4\,\text{m}$ rectangular workspace $\mathcal{W}^{\ast}$ cluttered with 8 circular obstacles, denoted as $\mathcal{O}_{i}^{\ast}$, whose centers and radii are listed in Table I. The distance between the virtual and actual control points is $\ell=0.05\,\rm m$, the radius of the circumscribed circle around $\bm{P}$ is $r=0.2\,\rm m$, and the clearance constant in Assumption 2 is $h=0.2\,\rm m$. The size of the safety margin is set to $\epsilon=0.1\,\rm m$, and the obstacle influence region is specified by $\epsilon^{\ast}=0.2\,\rm m$. The goal location is taken as $\mathbf{x}^{\ast}=[2.5,1]^{\top}\,\rm m$. Assume that the control input constraint is $u_{m}=1.5$, and the disturbance $\mathbf{u}_{d}$ is of the following form
$\mathbf{u}_{d}=0.01\times\left[\begin{matrix}\sin(0.2t)+1\\ \cos(0.3t)-2\end{matrix}\right].$ (47)
According to Remark 4, the planner parameters are chosen as $\alpha=0.03$ and $\beta=0.005$, while the controller parameters are chosen as $\rho=0.06$, $k=0.1$, $\psi=0.005$, $\eta=0.1$, $\gamma=0.01$, $d_{m}=0.03$, $\delta=0.005$, and $\hat{d}(0)=0.01$. The simulation duration is $500\,\rm s$, and the sample step is $0.01\,\rm s$. We begin by verifying the efficiency of the proposed tangent cone-based path planning method (denoted as “TC-based planner”) in the absence of disturbances. For comparison purposes, the PF-based path planning method presented in [30] (denoted as “PF-based planner”) is also simulated. A slight modification is made for [30] to facilitate its implementation. The attractive and repulsive potentials are chosen as
$U_{att}(\mathbf{x}_{d})=$ $\dfrac{1}{2}k_{a}\|\mathbf{x}_{d}-\mathbf{x}^{\ast}\|^{2},$
$U_{rep}(\mathbf{x}_{d})=$ $\dfrac{1}{2}k_{r}\left(\sum_{i=0}^{8}\dfrac{1}{\rho_{i}}\right)\|\mathbf{x}_{d}-\mathbf{x}^{\ast}\|^{2},$
where $k_{a}$ and $k_{r}$ are positive weights, $\rho_{0}:=1-(x_{d}/2.9)^{20}-(y_{d}/1.4)^{20}$, while $\rho_{i}:=\|\mathbf{x}_{d}-\mathbf{c}_{i}\|^{2}-(r+r_{i}+\epsilon)^{2}$, $i\in\mathbb{I}$. The total velocity is then given by
For fair comparison, we chose the parameters of the PF-based path planner to be $k_{a}=0.05$ and $k_{r}=0.0001$, such that the two planners have nearly identical convergence times. Figure 6 shows the obtained robot navigation trajectories of both path planners starting at a set of initial positions (purple points). As can be seen, both planners can successfully guide the virtual control point $\bm{P}$ from different initial positions to the goal (red point), while avoiding all obstacle regions $\mathcal{O}_{i}^{\epsilon}$, $i=1,2,...,8$. However, the proposed planner yields shorter trajectories than the PF-based one. The comparison results of the velocity norm $\|\bm{\tau}_{d}\|$ are depicted in Fig. 7, from which it is clear that the PF-based planner requires larger velocity than the proposed one. Actually, $\bm{\tau}_{d}$ under the proposed planner is upper bounded by $\alpha=0.03$, which is consistent with our analysis in Remark 3. The above results demonstrate the effectiveness and superiority of the proposed path planner.

### v2

@@ III-C Prescribed-time tube-following control
where $\rho>0$ is a user-defined constant. At each time instance, the actual position $\mathbf{x}$ of the virtual control point $\bm{P}$ is allowed to stay within a 1-sphere of radius $\rho$ and center $\mathbf{x}_{d}(t)$. Unifying all such spheres along $t$ forms a tube around $\mathbf{x}_{d}(t)$, as shown in Fig. 5. Since a safety margin is considered in the planning module, if the tube radius is taken no larger than the size $\epsilon$ of the safety margin and (22) holds for all $t\geq 0$, then the actual trajectory $\mathbf{x}(t)$ remains within the safe tube without colliding with any obstacles, that is, $\mathbf{x}(t)\in\mathcal{X}$, $\forall t\geq 0$.
Figure 5: Illustration of the predefined safe tube.
In the following, a tube-following controller is developed to achieve prescribed-time trajectory tracking, while ensuring that $\mathbf{x}(t)$ evolves within the tube defined by (22). To this end, we define a transformed error
@@
Inspecting (23) reveals that $10\leq\xi(t)<$ 1 is equivalent to (22), and $\xi(t)=0$ only when $\mathbf{x}_{e}(t)=\mathbf{0}$. Therefore, the prescribed-time tube-following control problem boils down to achieving $\lim_{t\to T_{f}}\xi(t)=0$, while ensuring $10\leq\xi(t)<$ 1, $\forall t\geq 0$, via a properly-designed controller.
To achieve trajectory tracking within a prescribed finite time $T_{f}>0$, as per Definition 2, we introduce a TSTF $t:=\eta_{f}(s)=T_{f}(1-e^{-s/T_{f}})$, which generates the following PTGF:
$\alpha_{f}(t)=\alpha_{f}(\eta_{f}(s)):=\dfrac{1}{\eta_{f}^{\prime}(s)}=\dfrac{T_{f}}{T_{f}-t},~{}t\in[0,T_{f}).$ (25)
Let us define $\mathbf{z}:=\mathbf{x}_{e}/(\rho^{2}(1-\xi))$ and design the prescribed-time control law as
$\mathbf{u}=\alpha_{f}^{\ast}(t)\mathbf{u}_{n}+\mathbf{R}^{-1}(\theta)(1-\alpha_{f}^{\ast}(t))\bm{\tau}_{d},$ (26)
@@
$\mathbf{u}_{n}=\mathbf{R}^{-1}(\theta)\left(-k_{1}\mathbf{x}_{e}-\dfrac{k_{2}}{\alpha_{f}^{\ast}(t)}\mathbf{z}+\bm{\tau}_{d}\right),$ (27)
where $k_{1},k_{2}>0$ are constant gains, and $\alpha_{f}^{\ast}(t)$ is a continuous PTGF akin to $\alpha^{\ast}(t)$ in (20), given by
$\alpha_{f}^{\ast}(t):=\left\{\begin{aligned} &\alpha_{f}(t),&~{}&t\in[0,T_{f}^{\ast}),\\ &\alpha_{f}(T_{f}^{\ast}),&~{}&t\in[T_{f}^{\ast},\infty),\end{aligned}\right.$ (28)
with $T_{f}^{\ast}:=T_{f}-\varsigma_{f}>0$ and $0<\varsigma_{f}\ll T_{f}$ being constants.
@@
Consider the WMR kinematics (3) under Assumption 1. If the initial position $\mathbf{x}(0)$ satisfies (22) and the gain $k_{1}$ is chosen such that $k_{1}T_{f}\gg 1$ holds, then the control law (26) achieves practical prescribed-time convergence of the position error $\mathbf{x}_{e}$, while ensuring it always remains within the safe tube defined by (22).
Proof. We analyze the closed-loop behavior over two time intervals $[0,T_{f}^{\ast})$ and $[T_{f}^{\ast},\infty)$, separately.
1) We first consider the time interval $t\in[0,T_{f}^{\ast})$. Substituting (26) into (21) yields
$\dot{\mathbf{x}}_{e}=-k_{1}\alpha_{f}(t)\mathbf{x}_{e}-k_{2}\mathbf{z}+\mathbf{R}(\theta)\mathbf{u}_{d}.$ (29)
To proceed, let us define
$\bar{\mathbf{x}}_{e}(s)$ $:=\mathbf{x}_{e}(\eta_{f}(s)),$ (30a)
$\bar{\xi}(s)$ $:=\xi(\eta_{f}(s))=\|\bar{\mathbf{x}}_{e}(s)\|^{2}/\rho^{2},$ (30b)
$\bar{\mathbf{z}}(s)$ $:=\mathbf{z}(\eta_{f}(s))=\bar{\mathbf{x}}_{e}(s)/(\rho^{2}(1-\bar{\xi}(s))).$ (30c)
Further define $\bar{\mathbf{d}}(s):=\mathbf{R}(\theta(\eta_{f}(s)))\mathbf{u}_{d}(\eta_{f}(s))$ for notational concision. Then, the closed-loop system (29) can be rewritten in the stretched infinite time interval $s\in[0,\infty)$ as
$\bar{\mathbf{x}}_{e}^{\prime}(s)$ $=\dfrac{d\bar{\mathbf{x}}_{e}(s)}{ds}=\dfrac{d\mathbf{x}_{e}(\eta_{f}(s))}{d\eta_{f}(s)}\cdot\dfrac{d\eta_{f}(s)}{ds}$
$=-k_{1}\bar{\mathbf{x}}_{e}(s)-\dfrac{1}{\alpha_{f}(\eta_{f}(s))}(k_{2}\bar{\mathbf{z}}(s)-\bar{\mathbf{d}}(s)),$ (31)
with $\bar{\mathbf{x}}_{e}(0)=\mathbf{x}_{e}(0)$ and $\bar{\mathbf{x}}_{e}^{\prime}(0)=\dot{\mathbf{x}}_{e}(0)$.
Consider the following logarithmic barrier function
$L(\bar{\xi})=\dfrac{1}{2}\ln\dfrac{1}{1-\bar{\xi}}.$ (32)
Taking the derivative of $L$ w.r.t. $s$ along (III-C) yields
$L^{\prime}(\bar{\xi}(s))=\bar{\mathbf{z}}^{\top}(s)\left(-k_{1}\bar{\mathbf{x}}_{e}(s)-\dfrac{k_{2}\bar{\mathbf{z}}(s)-\bar{\mathbf{d}}(s)}{\alpha_{f}(\eta_{f}(s))}\right).$ (33)
From [31, Lemma 3], it follows that $-k_{1}\bar{\mathbf{z}}^{\top}\bar{\mathbf{x}}_{e}=-k_{1}\bar{\xi}/(1-\bar{\xi})\leq-k_{1}L$ holds for all $10\leq\bar{\xi}<$ 1. In addition, recalling Assumption 1 and the fact that $\|\mathbf{R}(\theta)\|=1$, we have $\bar{\mathbf{z}}^{\top}\bar{\mathbf{d}}\leq k_{2}\|\bar{\mathbf{z}}\|^{2}+d/(4k_{2})$. As a result, (33) reduces to
$L^{\prime}(\bar{\xi}(s))$ $\leq-k_{1}L(\bar{\xi}(s))+\dfrac{d}{4k_{2}}\dfrac{T_{f}-\eta_{f}(s)}{T_{f}}$
$=-k_{1}L(\bar{\xi}(s))+\dfrac{d}{4k_{2}}e^{-\frac{s}{T_{f}}},$ (34)
on the set $10\leq\bar{\xi}<$ 1. Integrating both sides of (III-C) leads to
$L(\bar{\xi}(s))$ $\leq L(\bar{\xi}(0))e^{-k_{1}s}+De^{-\frac{s}{T_{f}}},$ (35)
where $D:=T_{f}d/(4k_{2}(k_{1}T_{f}-1))>0$ is a constant. From (35), it is clear that $L(\bar{\xi})$ is bounded, indicating that $10\leq\bar{\xi}(s)<$ 1 for all $s\geq 0$. As a result, $\|\bar{\mathbf{x}}_{e}(s)\|<\rho$ holds for all $s\geq 0$. Since $t\to T_{f}$ as $s\to\infty$, it follows from (30a) that $\mathbf{x}_{e}(t)$ keeps within the safe tube (22) for all $t\in[0,T_{f}^{\ast})$.
In the following, we show that $\mathbf{x}_{e}(t)$ converges to a small neighborhood of origin at $t=T_{f}^{\ast}$. It follows from (35) that $\lim_{s\to s^{\ast}}L(\bar{\xi}(s))\leq L^{\ast}:=L(\bar{\xi}(0))e^{-k_{1}s^{\ast}}+De^{-s^{\ast}/T_{f}}$. As per the definition of $\eta_{f}(s)$, one can easily deduce that $s\to s^{\ast}:=-T_{f}\ln(\varsigma_{f}/T_{f})$ as $t\to T_{f}^{\ast}$. Since $0<\varsigma_{f}\ll T_{f}$, $s^{\ast}$ is a very large value, which renders $L^{\ast}$ very small. By (23) and (32), we get $\lim_{s\to s^{\ast}}\|\bar{\mathbf{x}}_{e}(s)\|\leq\rho\sqrt{1-e^{-2L^{\ast}}}$, which implies that $\bar{\mathbf{x}}_{e}$ converges to a small neighborhood of origin as $s\to s^{\ast}$. Since $t\to T_{f}^{\ast}$ as $s\to s^{\ast}$ and $\bar{\mathbf{x}}_{e}(s)=\mathbf{x}_{e}(\eta_{f}(s))$ (see (30a)), it can be claimed that $\lim_{t\to T_{f}^{\ast}}\|\mathbf{x}_{e}(t)\|\leq\rho\sqrt{1-e^{-2L^{\ast}}}$.
2) Consider the time interval $t\in[T_{f}^{\ast},\infty)$, in which $\alpha_{f}^{\ast}(t)=\alpha_{f}(T_{f}^{\ast})\equiv T_{f}/\varsigma_{f}$ and the closed-loop system becomes
$\dot{\mathbf{x}}_{e}=-k_{1}\alpha_{f}(T_{f}^{\ast})\mathbf{x}_{e}-k_{2}\mathbf{z}+\mathbf{R}(\theta)\mathbf{u}_{d}.$ (36)
We likewise consider the barrier function $L$ in (32). Taking its derivative w.r.t. $t$ along (36) gets
$\dot{L}(\xi(t))\leq-\bar{k}L(\xi(t))+\bar{d},$ (37)
where $\bar{k}:=k_{1}T_{f}/\varsigma_{f}$ and $\bar{d}:=d/(4k_{2})$ are defined for notational concision. Solving (37), we get
$L(\xi(t))\leq L(\xi(T_{f}^{\ast}))e^{-\bar{k}(t-T_{f}^{\ast})}+\bar{d}/\bar{k},~{}\forall t\geq T_{f}^{\ast},$ (38)
indicating that $L(\xi(t))$ is bounded on $t\in[T_{f}^{\ast},\infty)$. Thus, $\mathbf{x}_{e}(t)$ satisfies the tube constraint (22) for all $t\geq T_{f}^{\ast}$. In addition, $L$ remains within the set $L(\xi(t))\leq L^{\ast}+\bar{d}/\bar{k}$ for all $t\geq T_{f}^{\ast}$. As a result, the tracking error $\mathbf{x}_{e}$ remains within the residual set $\|\mathbf{x}_{e}(t)\|\leq\rho\sqrt{1-e^{-2(L^{\ast}+\bar{d}/\bar{k})}}$ for all $t\geq T_{f}^{\ast}$.
@@
Under the InPTC framework, if the radius of the tube is set less than the safety margin $\epsilon$, whilst $T_{f}$ is set no larger than $T$, then the actual trajectory of $\mathbf{x}$ will converge to a small neighborhood of the goal position $\mathbf{x}^{\ast}$ (if no local minima occurs) within the prescribed time $T$, while avoiding collision with all obstacles $\mathcal{O}_{i}$, $i\in\mathbb{I}$ along the way. Therefore, the proposed InPTC can achieve prescribed-time collision-free navigation of WMRs. Furthermore, inspecting (35) and (38) reveals that choosing a smaller $\varsigma_{f}$ together with larger values for $k_{1}$ and $k_{2}$ decreases the size of residual set. However, this comes at the price of higher control gains (see (26)), particularly when selecting a very small $\varsigma_{f}$. This, in turn, may excite the unmodeled high frequency dynamics, leading to instability of the closed system. Thus, $k_{1}$, $k_{2}$, and $\varsigma_{f}$ should be judiciously chosen to strike a balance between tracking accuracy and system stability.
@@ IV Simulation Results
In this section, we demonstrate the proposed InPTC scheme in a $6.4\,\text{m}\times 3.4\,\text{m}$ rectangular workspace $\mathcal{W}^{\ast}$ cluttered with 8 circular obstacles, denoted as $\mathcal{O}_{i}^{\ast}$, whose centers and radii are listed in Table I. The distance between the virtual and actual control points is $\ell=0.05\,\rm m$, the radius of the circumscribed circle around $\bm{P}$ is $r=0.2\,\rm m$. The safety margin and influence region of obstacles are set to $\epsilon=0.1\,\rm m$ and $\epsilon^{\ast}=0.2\,\rm m$, respectively. The desired goal is $\mathbf{x}^{\ast}=[2.5,1]^{\top}\,\rm m$, and the task completion time is $200\,\rm s$. The disturbance $\mathbf{u}_{d}$ is modeled as $\mathbf{u}_{d}=0.01[\sin(0.2t)+1,\cos(0.3t)-2]^{\top}$. The simulation duration is $1000\,\rm s$, and the sample step is $0.05\,\rm s$.